The sum of two numbers is 21. Three times the smaller number is 2 less than twice the larger number. Find the two numbers
Answer by stanbon(48558)

The sum of two numbers is 21. Three times the smaller number is 2 less than twice the larger number. Find the two numbers
--------------
Let one of the numbers be x; the other number is 21-x.
EQUATION:
3x = 2(21-x)-2
3x = 42-2x-2
5x=40
x=8 (smaller number)
21-x=13 (other number)
=============
Cjeers.
StamJ/